# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

### v3

@@ III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
Spatial information is particularly crucial for DVGL. To enhance spatial invariance, inspired by [shen2023mccg], we design a dual-path spatial invariance enhancement learning module to mine spatial information from different dimensions.
@@
$\mathbf{f}_{v+}^{hu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu}))+\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu})))\times\mathbf{f}_{v}^{hu}$ (5)
$\mathbf{f}_{v+}^{wu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu}))+\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu})))\times\mathbf{f}_{v}^{wu}.$ (6)
@@ III-A3 Multi-level Optimization
In the CDIS, the features $\mathbf{f}_{v++}^{u++}$ obtained from the SIFL are optimized using the mean squared error loss $\mathcal{L}_{MSE}$ [xia2024enhancing], which improves the stability and consistency of the learned features. Furthermore, to enhance structural representation in the spatial invariance enhancement learning module, the features $\mathbf{f}_{v}^{au}$ are first extracted by the backbone and processed via global average pooling. The features $\mathbf{f}_{v}^{au}$ combined with $\mathbf{f}_{v+}^{hu}$ and $\mathbf{f}_{v+}^{wu}$, and the resulting combination are passed through a classifier module [shen2023mccg] to obtain $\mathbf{f}_{v+}^{cu}$. Subsequently, $\mathbf{f}_{v+}^{cu}$ are optimized with cross-entropy loss $\mathcal{L}_{CE}$ to enhance feature discrimination and cross-view consistency. In addition, to further enhance the feature representations, $\mathbf{f}_{v}^{au}$ undergo independent optimization using InfoNCE loss [oord2018representation] $\mathcal{L}_{InfoNCE}$. The final objective is to minimize the sum of all losses. It can be formulated as
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the coefficients to control the relative importance of the loss terms.
@@
The CDIS lies in dual optimization of structural and spatial invariance. On one hand, it integrates local and global structural information along the channel dimension and is guided by $\mathcal{L}_{MSE}$ to effectively constrain the feature search space, which enhances the structural representations. On the other hand, CDIS retains the original spatial dimensions and performs joint interaction learning along C-H and C-W, where $\mathcal{L}_{CE}$ is employed to further mine scale-aware contextual cues. This structure–spatial collaborative modeling enables the CDIS to achieve rapid convergence and learn highly discriminative representations, even with limited paired data.
@@ III-B Cross-Domain Transfer Sub-network
The CDTS further explores cross-view correlations based on the enhanced consistency and compactness established by the CDIS. Specifically, the CDTS incorporates dual-path contrastive learning tailored explicitly for DVGL.
@@ III-B1 Cross-View Feature Memory Initialization
As shown in Fig. 3, at the beginning of each training epoch, we initialize cross-view feature instances using the invariant knowledge learned by the CDIS and perform clustering to generate pseudo-labels. This process aims to mitigate the issue of low-quality pseudo-label initialization caused by suboptimal feature extraction. Subsequently, we construct cluster memories for each view, specifically for drone and satellite perspectives denoted as $\phi_{k}^{d}$ and $\phi_{l}^{s}$, by computing the average feature representations of drone and satellite instances within each cluster. This process can be formulated as
@@
where $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ are drone and satellite instance features, respectively. $i=1,\cdots,N$, $j=1,\cdots,M$. $\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{l}^{s}$ denote the $k$-th cluster set in drone and $l$-th cluster satellite view, respectively. $|\cdot|$ indicates the number of instances per cluster.
@@
In selecting the clustering algorithm, we refer to several unsupervised cross-domain retrieval works [dai2022cluster, adca] and adopt DBSCAN [ester1996density] for clustering. Unlike partition-based methods such as K-means [likas2003global], DBSCAN utilizes a density-based method, which handles clusters of arbitrary shapes and noisy data. It does not require a predefined number of clusters, nor does it force all points to be assigned to a cluster. Additionally, DBSCAN can identify and exclude noise points, which ensures the selection of valid samples even in cross-view challenges
@@ III-B2 Cross-View Feature Memory Updating
where $\alpha$ is momentum updating factor and $\omega$ is iteration number. $q_{d}$ and $q_{s}$ represent the given query features of the drone and the satellite, respectively.
@@ III-B3 Cross-View Contrastive Loss
$\mathcal{L}^{d}=-\log\frac{\exp(q_{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q_{d}\cdot\phi^{d}_{k}/\tau)}$ (11)
$\mathcal{L}^{s}=-\log\frac{\exp(q_{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{l=0}^{L}\exp(q_{s}\cdot\phi^{s}_{l}/\tau)}$ (12)
where $\tau$ denotes the temperature hyper-parameter. Drone cluster memories $\phi^{d}_{k}$ and satellite cluster memories $\phi^{s}_{l}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q_{d(s)}$ and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ correspond to the positive cluster memory associated with $q_{d}$ and $q_{s}$, respectively.
@@
The CDTS is jointly optimized with $\mathcal{L}^{d}$ and $\mathcal{L}^{s}$. Thus, the final optimization for CDIKTNet is denoted as $\mathcal{L}\!=\!\mathcal{L}_{1}\!+\!\mathcal{L}^{d}\!+\!\mathcal{L}^{s}$.
@@
i MuSe-Net[wang2024multiple] 100 74.48 77.83 88.02 75.10
TransFG[zhao2024transfg] 100 84.01 86.31 90.16 84.61
IFSs[ge2024multibranch] 100 86.06 88.08 91.44 85.73
MCCG[shen2023mccg] 100 89.40 91.07 95.01 89.93
Sample4Geo[deuser2023sample4geo] 100 92.65 93.81 95.14 91.39
CAMP[wu2024camp] 100 94.46 95.38 96.15 92.72
DAC[xia2024enhancing] 100 94.67 95.50 96.43 93.79
@@
i IFSs [ge2024multibranch] 100 77.57 81.30 89.50 91.40 92.58 94.21 97.40 97.92 93.75 79.49 97.50 90.52 97.50 96.03 100.00 97.66
MCCG[shen2023mccg] 100 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
Sample4Geo[deuser2023sample4geo] 100 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36 97.50 93.63 98.75 96.70 98.75 98.28 98.75 98.05
DAC[xia2024enhancing] 100 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
CAMP[wu2024camp] 100 95.40 96.38 97.63 98.16 98.05 98.45 99.33 99.46 96.25 93.69 97.50 96.76 98.75 98.10 100.00 98.85
@@
IV Experimental Results
IV-A Datasets and Evaluation Protocol
We evaluate CDIKTNet on two challenging benchmarks for DVGL. University-1652 dataset [zheng2020university] consists of drone, satellite, and ground-view images from 1,652 locations in 72 universities. Its training set includes 701 buildings from 33 universities, while the test set comprises 951 buildings from 39 universities. SUES-200 dataset [zhu2023sues] comprises 200 locations, which are split into 120 for training and 80 for testing. Each location provides a satellite image and drone images taken at four different altitudes, which encompass diverse environments such as parks, lakes, and buildings. We employ Recall@K (R@K) and Average Precision (AP) as evaluation metrics. Additionally, Ground Truth Ratio (GT ratio) denotes the proportion of paired data within the entire experimental dataset.
@@ IV-A Datasets and Evaluation Protocol
i MCCG[shen2023mccg] 100 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[deuser2023sample4geo] 100 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
DAC[xia2024enhancing] 100 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
CAMP[wu2024camp] 100 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
@@
UM AnyLoc[keetha2023anyloc] 0 42.08 57.92 69.19 -
SegVLAD[garg2024revisit] 0 68.19 77.32 77.18 -
@@
UM AnyLoc[keetha2023anyloc] 0 34.00 46.74 41.00 51.94
SegVLAD[garg2024revisit] 0 69.50 72.98 73.00 74.26
@@
TABLE VI: Comparisons between existing state-of-the-art unsupervised methods and CDIKTNet trained in setting $i$ in generalization capacity.
@@
UM AnyLoc[keetha2023anyloc] 0 30.00 42.74 34.00 46.74 36.78 50.41 41.00 51.97
SegVLAD[garg2024revisit] 0 56.00 59.08 69.50 72.98 72.50 74.10 73.00 74.26
@@
- Sample4Geo[deuser2023sample4geo] 2 62.10 68.11 69.85 74.93 74.85 79.14 76.40 80.33 77.50 57.91 88.75 70.17 91.25 76.99 90.00 78.89
CAMP[wu2024camp] 2 54.73 61.70 62.30 68.39 65.83 71.42 68.40 73.58 68.75 49.25 82.50 60.49 87.50 68.79 91.25 74.43
DAC[xia2024enhancing] 2 55.83 62.16 68.50 73.25 73.38 77.34 74.70 78.64 70.00 51.09 85.00 61.59 90.00 70.53 91.25 74.63
@@
- Sample4Geo[deuser2023sample4geo] 10 75.43 79.21 84.20 86.70 87.18 89.40 89.63 91.61 91.25 76.89 92.50 82.76 96.25 86.60 96.25 89.23
CAMP[wu2024camp] 10 76.27 80.39 83.83 86.89 89.66 91.62 90.85 92.71 92.50 77.50 96.25 86.63 97.50 91.45 97.50 93.29
DAC[xia2024enhancing] 10 76.80 80.65 85.05 87.65 88.80 90.81 90.58 92.32 93.75 79.06 97.50 86.78 97.50 90.58 98.75 92.54
@@
IV-B Implementation Details
During training and testing, all input images are resized to $3\times 384\times 384$. The CDIS is trained for 1 epoch, while the CDTS runs for 30 epochs, with a batch size of 64. In DBSCAN, the neighborhood radius is set to 0.40 and 0.30 for drone-view images and satellite-view images, respectively, and the minimum number required to form a cluster is 4. AdamW optimizer is used for the CDIS with an initial learning rate of 0.001, followed by SGD optimization for the CDTS with a learning rate of 0.00025. Furthermore, in Eq.(7), $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.6, 0.1, and 1, respectively. The CDTS is activated only when there exists unpaired data.
IV-C Comparison with State-of-the-art Methods
We conduct experiments with three settings, namely (i) fully supervised learning, (ii) few-shot initialization, and (iii) cross-domain initialization. The methods under comparison encompass a range of state-of-the-art techniques. For supervised learning, these include MuSe-Net [wang2024multiple], TransFG [zhao2024transfg], IFSs [ge2024multibranch], MCCG [shen2023mccg], Sample4Geo [deuser2023sample4geo], CAMP [wu2024camp] and DAC [xia2024enhancing]. In the realm of unsupervised learning, the methods are EM-CVGL [li2024learning], SegVLAD [garg2024revisit], AnyLoc [keetha2023anyloc] and Li [11010141]. Notably, SegVLAD [garg2024revisit], AnyLoc [keetha2023anyloc] and Li [11010141] are reported in separate Tables IV, V and VI due to the absence of complete results in their publications.
[…]
@@ IV-C Comparison with State-of-the-art Methods
- Sample4Geo[deuser2023sample4geo] 2 62.41 66.85 84.45 60.52
CAMP[wu2024camp] 2 62.53 67,01 83.31 59.26
DAC[xia2024enhancing] 2 63.68 68.16 87.59 61.26
@@
- CAMP[wu2024camp] 10 83.64 86.27 92.44 82.27
Sample4Geo[deuser2023sample4geo] 10 84.51 86.94 93.01 83.53
DAC[xia2024enhancing] 10 84.64 87.07 93.15 83.03
@@
(ii) Few-shot Initialization: To evaluate the effectiveness of CDIKTNet in leveraging unlabeled data for learning, we initialize CDIKTNet with only 2% and 10% of the paired data. The remaining training data are entirely unpaired. As shown in Tables I and II, CDIKTNet with only 2% of the paired data achieves performance comparable to or even surpassing that of IFSs [zhao2024transfg]. With 10% of paired data, CDIKTNet achieves further improvements, surpassing most fully supervised methods and achieving performance comparable to MCCG [shen2023mccg]. Notably, as shown in Table III, CDIKTNet in this setting trained with only 10% of the paired data exhibits even stronger generalization than these fully supervised methods.
@@
Furthermore, compared with state-of-the-art unsupervised methods, including EM-CVGL [li2024learning], AnyLoc [keetha2023anyloc], SegVLAD [garg2024revisit], and Li [11010141], as illustrated in Tables I, II, IV and V, CDIKTNet in this setting significantly outperforms them in all metrics on both datasets, despite relying on only a small proportion of paired data. In terms of generalization, CDIKTNet still significantly surpasses these methods, as reported in Tables III and VI. This underscores the significance of employing a limited quantity (e.g., 2%) of paired data for initialization, as it serves to steer the exploration toward revealing deeper cross-view representations from unpaired data.
@@
(iii) Cross-domain Initialization: We propose a novel experimental setting for cross-domain validation. We initialize CDIKTNet with 2% of paired data and then conduct the training process with the remaining unpaired data on one dataset that corresponds to a specific domain. This trained model is then transferred to another dataset representing a distinct domain for further training without requiring additional pairing relationships, with 0% of paired data during the transfer phase. Therefore, in this setting, the GT ratio is marked as 0%.
@@
The reverse validation follows the same procedure. As shown in Tables I, II, IV and V, after training in one domain, CDIKTNet can be directly transferred to another domain. It outperforms state-of-the-art unsupervised methods in all metrics and achieves performance comparable to several fully supervised methods. Moreover, as reported in Tables III and VI, CDIKTNet demonstrates remarkable generalization performance. It significantly outperforms these state-of-the-art unsupervised methods. Notably, as shown in V, the experimental results in this setting demonstrates superior performance compared with the few-shot initialization setting, which utilizes 2% of paired data. This indicates that strong domain adaptability enables the cross-domain initialization setting to outperform limited paired supervision under some cases.
@@
The (ii) few-shot initialization setting and (iii) cross-domain initialization setting are two typical scenarios of limited supervision. The former targets the current domain by utilizing a limited amount of paired data for initialization, which is a process of supervised learning. Subsequently, unpaired data is employed for unsupervised learning. The latter addresses the issue of a complete lack of paired data in a new domain. It first performs initialization through supervised learning in an existing domain using a limited amount of paired data. Then, it uses the remaining unpaired data for unsupervised learning. The obtained model is directly applied to unsupervised learning in the new domain, and has achieved remarkable results in the DVGL task in the new domain.

### v4

@@ III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
Spatial information is particularly crucial for DVGL. To enhance spatial invariance, inspired by [36], we design a dual-path spatial invariance enhancement learning module to mine spatial information from different dimensions.
@@
$\mathbf{f}_{v+}^{hu}$ $=\mathcal{F}\left(\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{hu}\right)\right)+\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{wu}\right)\right)\right)\times\mathbf{f}_{v}^{hu}$ (5)
$\mathbf{f}_{v+}^{wu}$ $=\mathcal{F}\left(\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{wu}\right)\right)+\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{hu}\right)\right)\right)\times\mathbf{f}_{v}^{wu}$ (6)
@@ III-A3 Multi-level Optimization
In the CDIS, the features $\mathbf{f}_{v++}^{u++}$ obtained from the SIFL, which denotes as $\mathbf{f}_{i++}^{d++}$ for drone-view intermediate features and $\mathbf{f}_{j++}^{s++}$ for satellite-view intermediate features, are optimized using the mean squared error loss $\mathcal{L}_{MSE}$ [48]. Furthermore, to enhance structural representation in the spatial invariance enhancement learning module, the features $\mathbf{f}_{i}^{ad}$ and $\mathbf{f}_{j}^{as}$ are obtained by applying global average pooling to $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ , respectively. The feature $\mathbf{f}_{i}^{ad}$ combined with $\mathbf{f}_{i+}^{hd}$ and $\mathbf{f}_{i+}^{wd}$ is passed through a classifier module [36] to obtain $\mathbf{f}_{i+}^{cd}$, as well as $\mathbf{f}_{j+}^{cs}$ obtained in the same manner. Subsequently, $\mathbf{f}_{i+}^{cd}$ and $\mathbf{f}_{j+}^{cs}$ are optimized with cross-entropy loss $\mathcal{L}_{CE}$ to enhance feature discrimination. In addition, to further enhance the feature representations, $\mathbf{f}_{i}^{ad}$ and $\mathbf{f}_{j}^{as}$ undergo independent optimization using InfoNCE loss $\mathcal{L}_{InfoNCE}$ [33]. The final objective is to minimize the sum of all losses. It can be formulated as
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ serve as loss coefficients to balance the relative contributions of individual loss terms.
@@
The CDIS lies in dual optimization of structural and spatial invariance. On one hand, it integrates local and global structural information along the channel dimension and is guided by $\mathcal{L}_{MSE}$ to effectively constrain the feature search space, which enhances the structural representations. On the other hand, the CDIS retains the original spatial dimensions and performs joint interaction learning along C-H and C-W, where $\mathcal{L}_{CE}$ is employed to further mine scale-aware contextual cues. This structure–spatial collaborative modeling enables CDIS to achieve rapid convergence and learn highly discriminative representations, even with limited paired data.
@@ III-B Cross-Domain Transfer Sub-network
The CDTS further explores cross-view correlations based on the enhanced compactness established by CDIS. Specifically, the CDTS incorporates dual-path contrastive learning tailored explicitly for DVGL.
@@ III-B1 Cross-View Feature Memory Initialization
As shown in Fig. 3, at the beginning of each training epoch, cross-view feature instances are initialized using the invariant knowledge learned by CDIS and then clustering is performed to generate pseudo-labels. It can mitigate the issue of low-quality pseudo-label initialization caused by suboptimal feature extraction. Subsequently, cluster memories are constructed for the drone-view and satellite-view. These memories $\{\phi_{k}^{d}\}_{k=1}^{K}$ and $\{\phi_{l}^{s}\}_{l=1}^{L}$ are obtained by computing the average feature representation of all drone or satellite instances within each cluster, where $K$ and $L$ denote the total number of drone-view clusters and satellite-view clusters, respectively. $\phi_{k}^{d}$ and $\phi_{l}^{s}$ serve as the cluster centroids for the drone-view and satellite-view features, respectively. This process can be formulated as
@@
where $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ are drone and satellite instance features, respectively. $i=1,\cdots,N$, $j=1,\cdots,M$. $\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{l}^{s}$ denote the $k$-th cluster in drone-view and $l$-th cluster in satellite-view, respectively. $|\cdot|$ indicates the number of instances per cluster.
@@
In selecting the clustering algorithm, we refer to several unsupervised cross-view retrieval works [[8, 51] and adopt DBSCAN [12] for clustering. Unlike partition-based methods such as K-means [26], DBSCAN utilizes a density-based method, which handles clusters of arbitrary shapes and noisy data. It does not require a predefined number of clusters, nor does it force all points to be assigned to a cluster. Additionally, DBSCAN can identify and exclude noise points, which ensures the selection of valid samples even in cross-view challenges.
@@ III-B2 Cross-View Feature Memory Updating
where $\alpha$ is the momentum factor and $\omega$ is the iteration number. $q_{d}$ and $q_{s}$ represent the given query features of the drone and the satellite, respectively.
@@ III-B3 Cross-View Contrastive Loss
$\mathcal{L}^{d}=-\log\frac{\exp(q_{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=1}^{K}\exp(q_{d}\cdot\phi^{d}_{k}/\tau)}$ (11)
$\mathcal{L}^{s}=-\log\frac{\exp(q_{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{l=1}^{L}\exp(q_{s}\cdot\phi^{s}_{l}/\tau)}$ (12)
where $\tau$ denotes the temperature hyper-parameter. $\phi^{d}_{k}$ (or $\phi^{s}_{l}$ ) is utilized as a feature vector at the cluster level to compute the distances between the query instance $q_{d}$ (or $q_{s}$) and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ correspond to the positive cluster memory associated with $q_{d}$ and $q_{s}$ in $\{\phi_{k}^{d}\}_{k=1}^{K}$ and $\{\phi_{l}^{s}\}_{l=1}^{L}$, respectively. In practice, for a given query feature, e.g., $q_{d}$ or $q_{s}$, the cluster memory that is assigned in the same cluster, namely $\phi^{d}_{+}$ or $\phi^{s}_{+}$, is designated as the positive sample.
@@
The CDTS is optimized jointly with $\mathcal{L}^{d}$ and $\mathcal{L}^{s}$. Thus, the final optimization for CDIKTNet is denoted as $\mathcal{L}\!=\!\mathcal{L}_{1}\!+\!\mathcal{L}^{d}\!+\!\mathcal{L}^{s}$.
IV Experimental Results
IV-A Datasets and Evaluation Protocol
CDIKTNet is evaluated on two challenging benchmarks for DVGL. University-1652 dataset [57] consists of drone, satellite, and ground-view images from 1,652 locations in 72 universities. Its training set includes 701 buildings from 33 universities, while the test set comprises 951 buildings from 39 universities. SUES-200 dataset [59] comprises 200 locations, which are split into 120 for training and 80 for testing. Each location provides a satellite-view image and drone-view images taken at four different altitudes, which encompass diverse environments such as parks, lakes, and buildings. We employ Recall@K (R@K) and Average Precision (AP) as evaluation metrics. Additionally, Ground Truth Ratio (GT ratio) denotes the proportion of paired data within the entire experimental dataset.
IV-B Implementation Details
During training and testing, all input images are resized to $3\times 384\times 384$. The CDIS is trained for 1 epoch, while the CDTS runs for 30 epochs, with a batch size of 64. In DBSCAN, the neighborhood radius is set to 0.40 and 0.30 for drone-view images and satellite-view images, respectively, and the minimum number required to form a cluster is 4. AdamW optimizer is used for CDIS with an initial learning rate of 0.001, followed by SGD optimization for the CDTS with a learning rate of 0.00025. Furthermore, in Eq.(7), $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.6, 0.1, and 1, respectively. The CDTS is activated only when there exist unpaired data.
IV-C Comparison with State-of-the-art Methods
We conduct experiments with three settings, namely (i) fully supervised learning, (ii) few-shot initialization, and (iii) cross-domain initialization. The methods in comparison encompass a range of state-of-the-art techniques. For supervised learning, these include MuSe-Net [45], TransFG [56], IFSs [15], MCCG [36], Sample4Geo [9], CAMP [47] and DAC [48]. In the realm of unsupervised learning, the methods are EM-CVGL [24], SegVLAD [14], AnyLoc [21] and Li [23]. Notably, SegVLAD [14], AnyLoc [21] and Li [23] are reported in separate Tables IV, V and VI due to the absence of complete results in their publications.
[…]
@@ IV-C Comparison with State-of-the-art Methods
i MuSe-Net[45] 100 74.48 77.83 88.02 75.10
TransFG[56] 100 84.01 86.31 90.16 84.61
IFSs[15] 100 86.06 88.08 91.44 85.73
MCCG[36] 100 89.40 91.07 95.01 89.93
Sample4Geo[9] 100 92.65 93.81 95.14 91.39
CAMP[47] 100 94.46 95.38 96.15 92.72
DAC[48] 100 94.67 95.50 96.43 93.79
@@
i IFSs [15] 100 77.57 81.30 89.50 91.40 92.58 94.21 97.40 97.92 93.75 79.49 97.50 90.52 97.50 96.03 100.00 97.66
MCCG[36] 100 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
Sample4Geo[9] 100 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36 97.50 93.63 98.75 96.70 98.75 98.28 98.75 98.05
DAC[48] 100 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
CAMP[47] 100 95.40 96.38 97.63 98.16 98.05 98.45 99.33 99.46 96.25 93.69 97.50 96.76 98.75 98.10 100.00 98.85
@@
i MCCG[36] 100 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[9] 100 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
DAC[48] 100 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
CAMP[47] 100 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
@@
UM AnyLoc[21] 0 42.08 57.92 69.19 -
SegVLAD[14] 0 68.19 77.32 77.18 -
@@
UM AnyLoc[21] 0 34.00 46.74 41.00 51.94
SegVLAD[14] 0 69.50 72.98 73.00 74.26
@@
TABLE VI: Comparisons between existing state-of-the-art unsupervised methods and CDIKTNet trained in setting ii in generalization capacity.
@@
UM AnyLoc[21] 0 30.00 42.74 34.00 46.74 36.78 50.41 41.00 51.97
SegVLAD[14] 0 56.00 59.08 69.50 72.98 72.50 74.10 73.00 74.26
@@
- Sample4Geo[9] 2 62.10 68.11 69.85 74.93 74.85 79.14 76.40 80.33 77.50 57.91 88.75 70.17 91.25 76.99 90.00 78.89
CAMP[47] 2 54.73 61.70 62.30 68.39 65.83 71.42 68.40 73.58 68.75 49.25 82.50 60.49 87.50 68.79 91.25 74.43
DAC[48] 2 55.83 62.16 68.50 73.25 73.38 77.34 74.70 78.64 70.00 51.09 85.00 61.59 90.00 70.53 91.25 74.63
@@
- Sample4Geo[9] 10 75.43 79.21 84.20 86.70 87.18 89.40 89.63 91.61 91.25 76.89 92.50 82.76 96.25 86.60 96.25 89.23
CAMP[47] 10 76.27 80.39 83.83 86.89 89.66 91.62 90.85 92.71 92.50 77.50 96.25 86.63 97.50 91.45 97.50 93.29
DAC[48] 10 76.80 80.65 85.05 87.65 88.80 90.81 90.58 92.32 93.75 79.06 97.50 86.78 97.50 90.58 98.75 92.54
@@
- Sample4Geo[9] 2 62.41 66.85 84.45 60.52
CAMP[47] 2 62.53 67,01 83.31 59.26
DAC[48] 2 63.68 68.16 87.59 61.26
@@
- CAMP[47] 10 83.64 86.27 92.44 82.27
Sample4Geo[9] 10 84.51 86.94 93.01 83.53
DAC[48] 10 84.64 87.07 93.15 83.03
@@
(ii) Few-shot Initialization: To evaluate the effectiveness of CDIKTNet in leveraging unlabeled data for learning, we initialize CDIKTNet with only 2% and 10% of the paired data. The remaining training data are entirely unpaired. As shown in Tables I and II, CDIKTNet with only 2% of the paired data achieves performance comparable to or even surpassing that of IFSs [56]. With 10% of paired data, CDIKTNet achieves further improvements, surpassing most fully supervised methods and achieving performance comparable to MCCG [36]. Notably, as shown in Table III, CDIKTNet in this setting trained with only 10% of the paired data exhibits even stronger generalization than these fully supervised methods.
@@
Furthermore, compared with state-of-the-art unsupervised methods, including EM-CVGL [24], AnyLoc [21], SegVLAD [14], and Li [23], as illustrated in Tables I, II, IV and V, CDIKTNet in this setting significantly outperforms them in all metrics on both datasets, despite relying on only a small proportion of paired data. In terms of generalization, CDIKTNet still significantly surpasses these methods, as reported in Tables III and VI. This underscores the significance of employing a limited quantity (e.g., 2%) of paired data for initialization, as it serves to steer the exploration toward revealing deeper cross-view representations from unpaired data.
@@
(iii) Cross-domain Initialization: A novel experimental setting is proposed for cross-domain validation. CDIKTNet is initialized with 2% of paired data on one dateset and then conducts the training process with the remaining unpaired data, whose whole samples that corresponds to a specific domain. This trained model is then transferred to another dataset representing a distinct domain for further training without requiring additional pairing relationships. Notably, this transfer learning phase operates with 0% of paired data. Therefore, in this setting, the GT ratio is marked as 0%.
In setting iii of Tables I, and IV, CDIKTNet is first initialized by using 2% paired data on SUES-200 dataset and then continue training with the remaining unpaired data. The resulting model has corresponded to setting ii with GT Ratio 2% in Table II. This model is subsequently used as a cross-domain initialization and transferred to the other dataset University-1652 for further training, where no paired data are required in the following process. Similarly, in setting iii of Tables II and V, CDIKTNet is initialized using 2% paired data on the University-1652 dataset and continue training with its remaining unpaired data. This trained model has reported in setting ii with GT Ratio 2% in Table I. This model is then used as a cross-domain initialization on the SUES-200 dataset, where the transfer stage also relies entirely on unpaired data on this new dataset. Furthermore, the results in setting iii of Tables III and VI are obtained by directly evaluating the models from setting iii of Tables I and IV on the SUES-200 dataset, without requiring any additional fine-tuning.
@@
As shown in Tables I, II, IV and V, after training in one domain, CDIKTNet can be directly transferred to another domain. It outperforms state-of-the-art unsupervised methods in all metrics and achieves performance comparable to several fully supervised methods. Moreover, as reported in Tables III and VI, CDIKTNet demonstrates remarkable generalization performance. It significantly outperforms these state-of-the-art unsupervised methods. Notably, as shown in V, the experimental results in this setting demonstrates superior performance compared with the few-shot initialization setting, which has utilized 2% of paired data in current domain. This indicates that strong domain adaptability enables the cross-domain initialization setting to outperform limited paired supervision under some cases.
@@
The (ii) few-shot initialization setting and (iii) cross-domain initialization setting are two typical scenarios of limited supervision. The former targets the current domain by using a limited amount of paired data for initialization, which is a process of supervised learning. Subsequently, unpaired data is employed for unsupervised learning. The latter addresses the issue of a complete lack of paired data in a new domain. First, it performs initialization through supervised learning in an existing domain using a limited amount of paired data. Then, it uses the remaining unpaired data for unsupervised learning. The obtained model is directly applied to unsupervised learning in the new domain, and has achieved remarkable results in the DVGL task in the new domain.